### v5

@@ 8. Related Work
More recent work called Trocq (Cohen et al., 2024) implements an external form of transfer for Coq that directly supports relations more general than equivalences, like CoqEAL, but also supports proofs. Like Pumpkin Pi, Trocq goes out of its way to avoid depending on axioms like univalence and functional extensionality. Trocq’s motivation of supporting transfer of proofs across relations more general than equivalences is similar to the motivation of our extensions to Pumpkin Pi, with two differences: (1) our work supports a more limited class of relations that can be described as equivalences between quotient types, and (2) for that class of changes, by extending Pumpkin Pi’s proof term transformation, our work makes it possible to remove the old version of a type after applying repair. The major benefit of our tool relative to Trocq comes from (2)—while all proof repair tools implement a kind of transfer, not all transfer methods implement repair, and Trocq does not appear to implement repair.
@@ 9. Conclusions & Future Work
We extended Pumpkin Pi to support changes represented by quotient type equivalences, adding more expressive power. The key challenge we overcame was supporting quotient types in a proof repair algorithm built for a type theory that does not even have quotient types to begin with. We addressed this by representing quotient types using setoids, extending the Pumpkin Pi algorithm and implementation to repair proofs about equivalence relations, and adding new automation to dispatch newly generated proof obligations. Our extension demonstrated success on three case studies not supported by the original Pumpkin Pi.
Going forward, we would like to reconcile our grounding in Cubical Agda with our implementation in Coq, ideally getting the best of both worlds—something we could do easily if only we knew of a univalent language with both native quotient types and a strong metalanguage for building automation. We also hope to continue to improve our extension’s automation and usability. And we hope to look at other kinds of types and relations that can be expressed even when the type theory lacks them, as quotient types can be by way of setoids. We hope all of this will open the door to proof repair for more sophisticated classes of changes.
Appendix A Types of theorems that repair of eliminators is correct
In Section 7.2, we discuss proving theorem stating the correctness condition from the Pumpkin Pi paper. In Figure 13, we show that the Lam rule is correct. In Figure 14, the type of our proof that repair of a dependent eliminator is correct is presented. This theorem shows that, if all the inputs to the eliminator correspond to each other across the isomorphism, then the output of the eliminator applications also corresponds across that isomorphism. The term witnessing this type can be found in equivalence_int_abs.agda in the supplementary material.

### v6

@@ 7. Related Work
More recent work called, Trocq (Cohen et al., 2024) implements external transfer for Coq that directly supports relations more general than equivalences, like CoqEAL, but also supports proofs. Like Pumpkin Pi, Trocq goes out of its way to avoid depending on axioms like univalence and functional extensionality. Trocq’s motivation of supporting transfer of proofs across relations more general than equivalences is similar to the motivation of our extensions to Pumpkin Pi, with two differences: (1) our work supports a more limited class of relations that can be described as equivalences between quotient types, and (2) for that class of changes, by extending Pumpkin Pi’s proof term transformation, our work makes it possible to remove the old version of a type after applying repair. The major benefit of our tool relative to Trocq comes from (2)—while all proof repair tools implement a kind of transfer, not all transfer methods implement repair, and Trocq does not implement repair.
@@ 8. Conclusions & Future Work
We extended Pumpkin Pi to support changes represented by quotient type equivalences, enabling repair in situations previously untenable. The key challenge we overcame was supporting quotient types in a proof repair algorithm built for a type theory that does not have quotient types to begin with. We addressed this by representing quotient types using setoids, extending the Pumpkin Pi algorithm and implementation to repair proofs about equivalence relations, and adding new automation to dispatch newly generated proof obligations. Our extension demonstrated success on three case studies not supported by the original Pumpkin Pi. We also constructed the first internal correctness proofs for repair. We wish to continue to improve our extension’s automation and usability, and we hope to look at other kinds of types and relations that can be expressed even when the type theory lacks them, as quotient types can be by way of setoids. We hope this will open the door to proof repair for more sophisticated classes of changes.